OpenStudy (anonymous):

A city bus collected $780 in fares on one day. The price of a regular fare was $0.80, and the price of a discount fare was $0.40. If a total of 1200 people paid the fares on this bus, how many people paid the regular fare? A 1000 B 1950 C 600 D 750

OpenStudy (mertsj):

x= number of regular fares 1200-x= number of discount fares .80x+.40(1200-x)=780

OpenStudy (mertsj):

Can you solve the equation?

OpenStudy (anonymous):

No, I'm still stuck

OpenStudy (mertsj):

Multiply every term by 100 to remove all the decimal points.

OpenStudy (anonymous):

I still don't get it

OpenStudy (mertsj):

80x+40(1200-x)=78000

OpenStudy (anonymous):

then do i plug in my answer into the equation or something?

OpenStudy (mertsj):

Remove the parentheses by using the distributive property

OpenStudy (mertsj):

80x+48000-40x=78000 40x+48000=78000 40x=30000 x=750 750 people paid the regular fare.

OpenStudy (anonymous):

Where is the 48000 from?

OpenStudy (mertsj):

40 times 12000

OpenStudy (anonymous):

Ohh i see. I get it now. Thank you!

OpenStudy (mertsj):

yw